Give that T(x)=-2x^2-5x-8, evaluate :T(2)
Found 2 solutions by Nate, Osran_Shri:
Answer by Nate(3500)

T(x) = -2x^2 - 5x - 8
T(2) = -2*2^2 - 5*2 - 8 = -8 - 10 - 8 = -26

Answer by Osran_Shri(18)

Here is the solution:
T(x)=-2x^2-5x-8
We can get T(2) by substituting x=2 in the above equation as below,
T(2) = -2*2^2 - 5*2 - 8
T(2) = -8-10-8 = -26
Thats all...
B.Shridhar